# Prioritized Information Bottleneck Theoretic Framework with Distributed Online Learning for Edge Video Analytics

### v2

@@ Abstract
Collaborative perception systems leverage multiple edge devices, such surveillance cameras or autonomous cars, to enhance sensing quality and eliminate blind spots. Despite their advantages, challenges such as limited channel capacity and data redundancy impede their effectiveness. To address these issues, we introduce the Prioritized Information Bottleneck (PIB) framework for edge video analytics. This framework prioritizes the shared data based on the signal-to-noise ratio (SNR) and camera coverage of the region of interest (RoI), reducing spatial-temporal data redundancy to transmit only essential information. This strategy avoids the need for video reconstruction at edge servers and maintains low latency. It leverages a deterministic information bottleneck method to extract compact, relevant features, balancing informativeness and communication costs. For high-dimensional data, we apply variational approximations for practical optimization. To reduce communication costs in fluctuating connections, we propose a gate mechanism based on distributed online learning (DOL) to filter out less informative messages and efficiently select edge servers. Moreover, we establish the asymptotic optimality of DOL by proving the sublinearity of their regrets. To validate the effectiveness of the PIB framework, we conduct real-world experiments on three types of edge devices with varied computing capabilities. Compared to five coding methods for image and video compression, PIB improves mean object detection accuracy (MODA) while reducing 17.8% and reduces communication costs by 82.65% under poor channel conditions.
@@ I-A Background
Video analytics is rapidly transforming various sectors such as urban planning, retail analysis, and autonomous navigation by converting visual data streams into useful insights[2]. A large number of video cameras produce vast amounts of video data continuously and often require real-time video stream[3]. Numerous developing applications such as remote patient care[4], video games[5], UAV sensing[6, 7] and virtual and augmented reality depend on the efficient analysis of video data with minimal delay[8].
@@
The increasing number of smart devices requires a computational paradigm shift towards edge computing. This approach involves processing data closer to its source, resulting in several benefits compared to traditional cloud-based paradigms, particularly reduced latency and bandwidth costs. Even a delay as short as second can lead to disastrous consequences. For example, interactive applications such as online gaming and video conferencing require latencies below 100 ms to ensure real-time feedback and seamless user experience[9]. Similarly, VR/AR applications demand extremely low latencies, often less than 20 ms, to prevent motion sickness and maintain a high-quality immersive experience[10]. Utilizing remote cloud services for data processing can result in significant latency increases, often exceeding 100 ms[11]. Moreover, the importance of privacy, particularly in regions with strict data protection laws such as the General Data Protection Regulation (GDPR), makes edge computing even more attractive[12]. According to the Ponemon Institute, 60% of companies express apprehension toward cloud security and decide to manage their own data onsites in order to mitigate potential risks[13].
@@
A key aspect of video analytics, particularly for Bird’s Eye View (BEV) applications, is accurately capturing pedestrian occupancy across multiple camera views[14, 15]. BEV representations rely on precise spatial context to depict ground-level scenes in order to minimize occlusions, blind spots, and viewpoint discrepancies. Pedestrian occupancy data enables reliable identification and localization of individuals within a shared view, refining collaborative perception and enhancing tasks such as detection and prediction in complex, dynamic environments. However, the integration of edge devices into video analytics also brings in many significant challenges[16]. The computational demands of deep neural network (DNN) models, such as GoogLeNet[17], which requires about 1.5 billion operations per image classification, place a substantial burden on the limited processing capacities of edge devices[18]. Additionally, the outputs from high-resolution cameras increase the communication load. For example, a 4K video stream requires up to 18 Gbps of bandwidth to transmit raw video data, potentially overwhelming wireless networks[19]. Therefore, we need to explore efficient video coding for compressing streamed videos. As shown in Fig. 1(a), the traditional compression is to reconstruct streaming frame through efficient entropy coding and motion prediction. However, there are still extensive less informative data being processed, wasting communication bandwidth. For instance, if the tasks involve human recognition or positioning, the reconstructed background of each frame might not be useful for the application.
@@
As shown in Fig. 1(b), the information bottleneck (IB) framework is a feasible choice for task-oriented video compression, enabling a trade-off between communication cost and prediction accuracy for specific tasks. However, the current communication strategies for integrating edge devices into video analytics are not effective enough. One major issue is how to handle the computational complexity and transmission of redundant data generated from the overlapping fields of view (FOVs) from multiple cameras[20]. In scenarios with dense camera deployments, up to 60% of data can be redundant due to overlapping FOVs, which unnecessarily overburdens the network[21]. In addition, these strategies often lack adaptability in transmitting tailored data features based on Region of Interest (RoI) and signal-to-noise ratio (SNR), resulting in poor video fusion or alignment. These limitations can negatively impact collaborative perception, even making it less effective than single-camera setups[14].
@@ I-B State-of-the-Art
This subsection reviews advancements in edge video analytics, with an emphasis on the designs on communication-computing latency reduction. We explore the information bottleneck method to enhance task-oriented performance by minimizing data redundancy. Additionally, we investigate online learning for dynamic ROI management and perceptual quality.
@@ I-B1 Edge Video Analytics
(a) Traditional compression method with redundant data
(b) Information bottleneck method for task-specific compression
@@
Live video analytics is crucial in various domains such as autonomous driving[14, 22, 23, 15, 24, 25], mixed reality[26, 27], 3D point cloud analytics[28], and traffic control[29]. These applications, including object recognition[30], are typically equipped with sophisticated machine learning models like Convolutional Neural Networks (CNNs)and Graph Neural Networks (GNNs)[31, 32]. However, offloading these applications to central clouds can result in unpredictable transmission delays in wide area networks, particularly when streaming high-quality videos[33]. Therefore, researchers utilize edge computing to serve as a promising alternative to reduce service latency and energy consumption. Li et al. propose a novel approach called ESMO to optimize frame scheduling and model caching for edge video analytics[34]. Khani et al. introduce RECL, a new framework for video analytics that integrates model reuse and online retraining to quickly adapt expert models to specific video scenes, optimizing resource allocation and achieving substantial performance gains over prior methods[35]. Wang et al. design an MEC-enabled multi-device video analytics system using a Markov decision process to address real-time ground truth absence and content-varying degradation-accuracy issues that significantly enhances the accuracy-latency tradeoff through adaptive information gathering and efficient bandwidth allocation. However, few works consider how to strike a dynamic balance between channel resources and inference performance in a multi-camera sensing system.
@@
In typical edge video analytics scenarios, the lack of infrastructure and limited bandwidth makes real-time object detection challenging, especially for the multi-view camera sensing for wild animals or criminals in remote areas[36]. To achieve real-time object sensing, it is crucial to reduce redundant information and the bandwidth resource demand. Semantic communication can address this challenge by transmitting only the essential semantic information, thereby compressing data streams and reducing transmission overhead. Zhang et al. propose a comprehensive framework to highlight the importance of semantic communication in optimizing information transmission[37]. Shao et al. introduce a new conceptualization of semantic communication that characterizes it within joint source-channel coding theory, aiming to minimize the semantic distortion-cost region[38]. Xie et al. explore a deep learning-based semantic communication system with memory, showing how dynamic transmission techniques can enhance transmission reliability and efficiency by masking unessential elements[39]. Zhou et al. design and implement a deep learning-based image processing pipeline on the ESP32-CAM, proposing a DRL-based approach for efficient camera configuration adaptation in multi-camera systems[40]. Existing research primarily focuses on rate-distortion (R-D) optimization, adapting the bitstream rate based on channel state information (CSI) to reconstruct raw videos[41]. However, these methods rarely consider the performance of specific downstream tasks, such as mean object detection accuracy (MODA), as a system evaluation metric. Consequently, the transmitted information often contains redundant data.
@@
To address this issue, researchers incorporate the information bottleneck (IB) framework to optimize edge video analytics by focusing on task-specific performance, thereby reducing redundancy[42]. The IB framework helps the cause in selectively transmitting only the most relevant features needed for specific tasks, enhancing efficiency. Pensia et al. propose a novel feature extraction strategy in supervised learning that enhances classifier robustness to small input perturbations by incorporating a Fisher information penalty into the information bottleneck framework[43]. Wang et al. present a deep multi-view subspace clustering framework to extend the information bottleneck principle to a self-supervised setting, leading to superior performance in multi-view subspace clustering on real-world datasets[44]. IB tradeoff is well-suited for bandwidth-limited edge inference and is a key design principle in our study for efficient communication. Wang et al. introduce the Informative Multi-Agent Communication (IMAC) method, which uses the information bottleneck principle to develop efficient communication protocols and scheduling for multi-agent reinforcement learning under limited bandwidth[45]. Shao et al. propose a task-oriented communication scheme for multi-device cooperative edge inference, optimizing local feature extraction and distributed feature encoding to minimize data redundancy and focus on task-relevant information, leveraging the information bottleneck principle and extending it to a distributed deterministic information bottleneck framework. However, these existing studies often neglect the need to prioritize different data from multiple cameras for various downstream tasks, such as considering ROIs. Moreover, most existing studies overlook the correlation between multiple cameras in multi-view scenarios. Shao et al. extract compact task-oriented representations based on the IB principle, but they overlook the fact that different tasks require varying levels of priority[16]. By leveraging these correlations and levels of priority, it is possible to further reduce data rates by minimizing duplicate information across different camera feeds and enhance inference performance at the same time.
@@ I-B2 Learning-Based Transmission Scheduling
Online learning in multi-agent deep reinforcement learning (MADRL) enhances multi-camera sensing under dynamic channels and overlapped ROIs. Effective transmission scheduling determines when and which agents communicate through binary vectors indicating allowed communications at specific time steps, forming a communication graph. Central transmission scheduling schemes use a globally shared policy to control communication. Kim et al. propose SchedNet, which uses a global scheduler to limit broadcasting agents and reduce communication overhead[46]. Du et al. introduce FlowComm, forming a directed graph for communication[47]. Liu et al. develop GA-Comm, using a two-stage attention network (G2ANet) to manage agent interactions[48]. Niu et al. present MAGIC, a framework using a directed communication graph for enhanced coordination[49]. In distributed transmission scheduling schemes, each agent individually determines whether to communicate, forming a graph structure. Liu et al. propose a framework for multi-agent collaborative perception, addressing communication group construction and decision-making for efficient bandwidth use, significantly reducing communication while maintaining performance[50]. Deep learning optimizes these systems by refining communication actions and schedules, transmitting only relevant information, and minimizing redundancy. However, existing multi-camera cooperative sensing algorithms do not effectively address the transmission scheduling problem, particularly under dynamic wireless channels and overlapped ROIs.
@@ I-C Our Contributions
  • We propose the PIB framework that prioritizes the share data based on the signal-to-noise ratio (SNR) and camera coverage of the region of interest (RoI), reducing redundancy both spatially and temporally. This approach avoids the need for video reconstruction at edge servers and maintains low latency.
@@
  • To reduce communication costs in fluctuating links, we introduce a gate mechanism based on distributed online learning (DOL) to filter out unprofitable messages and efficiently select edge servers. We establish the asymptotic optimality of DOL by showing the sublinearity of their regrets.
@@ II System Model
As illustrated in Fig. 2, our system comprises a set of edge cameras, denoted as $\mathcal{K}=\{1,2,\ldots,K\}$, and a set of edge servers, denoted as $\mathcal{S}=\{1,2,\ldots,S\}$. Each camera has a specific Field of View (FoV), $\text{FoV}_{k}$, covering a subset of the monitored area. Our goal is to enable collaborative perception for pedestrian occupancy prediction under constrained channel capacity. The backbone of our system employs intermediate collaboration, where only encoded feature representations $Z$ are transmitted to a central aggregate server for decoding and inference. Other edge servers serve solely as relay nodes, as they do not perform decoding or final inference due to the network structure while does not support task offloading for these operations. The aggregate server is chosen to minimize overall system delay, while the relay-only role of other edge servers ensures streamlined data flow. The key notations are given in Table I for the ease of reading.
@@ III Problem Formulation
In this section, we establish the theoretical foundation for our PIB framework. We begin by detailing the IB analysis to determine the optimal balance between data compression and relevant information retention. Following this, we formulate the combinatorial multi-armed band (CMAB) problem to model the decision-making process of cameras in a distributed environment.
@@ III-A Prioritized Information Bottleneck Analysis
III-B Combinatorial Multiarmed Bandit (CMAB) Problem
@@
where $\Delta M_{k}=M_{\mathcal{K}\cup\{k\}}-M_{\mathcal{K}}$, and $\mathcal{K}$ is the set of cameras selected at time $t$. The computational cost of multi-camera fusion and inference at the edge server $s$ is denoted as $\Psi^{F}_{s}$. The computational cost of simply forwarding features from one edge server to another is denoted as $\Psi^{T}_{s}$. The remaining computing capacity of the $k$-th edge server is $\Psi^{R}_{k}$. $\mathcal{T}_{k,s}^{c\rightarrow e}$ denotes the latency of the transmission between the $k$-th camera and the $s$-th edge server, and $\mathcal{T}_{s}^{e\rightarrow e}$ denotes the latency of the transmission between the $k$-th edge server and the $s$-th edge server. Therefore, the CMAB problem can be formulated as:
@@
where (11a) ensures the number of selected cameras falls within the specified range and $|\mathcal{K}|=\sum_{k\in\mathcal{K}}{\mathcal{E}_{i,s}^{c\rightarrow e}}$, (11b) ensures that the remaining computing capacity of the aggregate server ($s_{0}$) chosen for inference is no less than the required capacity for fusion, (11c) ensures that each camera uses a unique transmission connection, (11d) ensures that the number of connections established by a single edge server does not exceed the maximum allowable connections, and (11e) ensures that the total latency for any transmission path is within the allowable time limit $\mathcal{T}^{U}$ for all edge servers ${s}\in\mathcal{S}$.

### v3

@@ Abstract
Collaborative perception systems leverage multiple edge devices, such as surveillance cameras or autonomous cars, to enhance sensing quality and eliminate blind spots. Despite their advantages, challenges such as limited channel capacity and data redundancy impede their effectiveness. To address these issues, we introduce the Prioritized Information Bottleneck (PIB) framework for edge video analytics. This framework prioritizes the shared data based on the signal-to-noise ratio (SNR) and camera coverage of the region of interest (RoI), reducing spatial-temporal data redundancy to transmit only essential information. This strategy avoids the need for video reconstruction at edge servers and maintains low latency. It leverages a deterministic information bottleneck method to extract compact, relevant features, balancing informativeness and communication costs. For high-dimensional data, we apply variational approximations for practical optimization. To reduce communication costs in fluctuating connections, we propose a gate mechanism based on distributed online learning (DOL) to filter out less informative messages and efficiently select edge servers. Moreover, we establish the asymptotic optimality of DOL by proving the sublinearity of its regrets. To validate the effectiveness of the PIB framework, we conduct real-world experiments on three types of edge devices with varied computing capabilities. Compared to five coding methods for image and video compression, PIB improves mean object detection accuracy (MODA) by 17.8% while reducing communication costs by 82.65% under poor channel conditions.
@@ I-A Background
Video analytics is rapidly transforming various sectors such as urban planning, retail analysis, and autonomous navigation by converting visual data streams into useful insights[2]. A large number of video cameras produce vast amounts of video data continuously and often require real-time video streams[3]. Numerous developing applications such as remote patient care[4], video games[5], UAV sensing[6, 7] and virtual and augmented reality depend on the efficient analysis of video data with minimal delay[8].
@@
The increasing number of smart devices requires a computational paradigm shift towards edge computing. This approach involves processing data closer to its source, resulting in several benefits compared to traditional cloud-based paradigms, particularly reduced latency and bandwidth costs. Even a delay as short as one second can have disastrous consequences. For example, interactive applications such as online gaming and video conferencing require latencies below 100 ms to ensure real-time feedback and a seamless user experience[9]. Similarly, VR/AR applications demand extremely low latencies, often less than 20 ms, to prevent motion sickness and maintain a high-quality immersive experience[10]. Utilizing remote cloud services for data processing can result in significant latency increases, often exceeding 100 ms[11]. Moreover, the importance of privacy, particularly in regions with strict data protection laws such as the General Data Protection Regulation (GDPR), makes edge computing even more attractive[12]. According to the Ponemon Institute, 60% of the companies express apprehension towards cloud security and decide to manage their own data on-site to mitigate potential risks[13].
@@
A key aspect of video analytics, particularly for Bird’s Eye View (BEV) applications, is accurately capturing pedestrian occupancy across multiple camera views[14, 15]. BEV representations rely on precise spatial context to depict ground-level scenes in order to minimize occlusions, blind spots, and viewpoint discrepancies. Pedestrian occupancy data enables reliable identification and localization of individuals within a shared view, refining collaborative perception and enhancing tasks such as detection and prediction in complex, dynamic environments. However, the integration of edge devices into video analytics also brings in many significant challenges[16]. The computational demands of deep neural network (DNN) models, such as GoogLeNet[17], which requires about 1.5 billion operations per image classification, place a substantial burden on the limited processing capacities of edge devices[18]. Additionally, the outputs from high-resolution cameras increase the communication load. For example, a 4K video stream requires up to 18 Gbps of bandwidth to transmit raw video data, potentially overwhelming wireless networks[19]. Therefore, we need to explore efficient video coding for compressing streamed videos. As shown in Fig. 1(a), traditional compression methods reconstruct streaming frames through efficient entropy coding and motion prediction. However, there is still a large amount of less informative data being processed, wasting communication bandwidth. For instance, if the tasks involve human recognition or positioning, the reconstructed background of each frame might not be useful for the application.
@@
As shown in Fig. 1(b), the information bottleneck (IB) framework is a feasible choice for task-oriented video compression, enabling a trade-off between communication cost and prediction accuracy for specific tasks. However, the current communication strategies for integrating edge devices into video analytics are not effective enough. One major issue is how to handle the computational complexity and transmission of redundant data generated from the overlapping fields of view (FOVs) of multiple cameras[20]. In scenarios with dense camera deployments, up to 60% of data can be redundant due to overlapping FOVs, which unnecessarily overburdens the network[21]. In addition, these strategies often lack adaptability in transmitting tailored data features based on Region of Interest (RoI) and signal-to-noise ratio (SNR), resulting in poor video fusion or alignment. These limitations can negatively impact collaborative perception, even making it less effective than single-camera setups[14].
@@ I-B State-of-the-Art
This subsection reviews advancements in edge video analytics, with an emphasis on the designs for communication-computing latency reduction. We explore the information bottleneck method to enhance task-oriented performance by minimizing data redundancy. Additionally, we investigate online learning for dynamic ROI management and perceptual quality.
@@ I-B1 Edge Video Analytics
(a) Traditional compression method with redundant data.
(b) Information bottleneck method for task-specific compression.
@@
Live video analytics is crucial in various domains such as autonomous driving[14, 22, 23, 15, 24, 25], mixed reality[26, 27], 3D point cloud analytics[28], and traffic control[29]. These applications, including object recognition tasks[30], are typically equipped with sophisticated machine learning models like Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs)[31, 32]. However, offloading these applications to centralized clouds can result in unpredictable transmission delays in wide area networks, particularly when streaming high-quality videos[33]. Therefore, researchers utilize edge computing to serve as a promising alternative to reduce service latency and energy consumption. Li et al. propose a novel approach called ESMO to optimize frame scheduling and model caching for edge video analytics[34]. Khani et al. introduce RECL, a new framework for video analytics that integrates model reuse and online retraining to quickly adapt expert models to specific video scenes, optimizing resource allocation and achieving substantial performance gains over prior methods[35]. Wang et al. design an MEC-enabled multi-device video analytics system using a Markov decision process to address real-time ground truth absence and content-varying degradation-accuracy issues, significantly enhancing the accuracy-latency tradeoff through adaptive information gathering and efficient bandwidth allocation. However, few works consider how to strike a dynamic balance between channel resources and inference performance in a multi-camera sensing system.
@@
In typical edge video analytics scenarios, the lack of infrastructure and limited bandwidth makes real-time object detection challenging, especially for the multi-view camera sensing for wildlife monitoring or security surveillance in remote areas[36]. To achieve real-time object sensing, it is crucial to reduce redundant information and the bandwidth demand. Semantic communication can address this challenge by transmitting only the essential semantic information, thereby compressing data streams and reducing transmission overhead. Zhang et al. propose a comprehensive framework to highlight the importance of semantic communication in optimizing information transmission[37]. Shao et al. introduce a new conceptualization of semantic communication that characterizes it within joint source-channel coding theory, aiming to minimize the semantic distortion-cost region[38]. Xie et al. explore a deep learning-based semantic communication system with memory, showing how dynamic transmission techniques can enhance transmission reliability and efficiency by masking unessential elements[39]. Zhou et al. design and implement a deep learning-based image processing pipeline on the ESP32-CAM, proposing a DRL-based approach for efficient camera configuration adaptation in multi-camera systems[40]. Existing research primarily focuses on rate-distortion (R-D) optimization, adapting the bitstream rate based on channel state information (CSI) to reconstruct raw videos[41]. However, these methods rarely consider the performance of specific downstream tasks, such as mean object detection accuracy (MODA), as a system evaluation metric. Consequently, the transmitted information often contains redundant data.
@@
To address this issue, researchers incorporate the information bottleneck (IB) framework to optimize edge video analytics by focusing on task-specific performance, thereby reducing redundancy[42]. The IB framework helps in selectively transmitting only the most relevant features needed for specific tasks, enhancing efficiency. Pensia et al. propose a novel feature extraction strategy in supervised learning that enhances classifier robustness to small input perturbations by incorporating a Fisher information penalty into the information bottleneck framework[43]. Wang et al. present a deep multi-view subspace clustering framework to extend the information bottleneck principle to a self-supervised setting, leading to superior performance in multi-view subspace clustering on real-world datasets[44]. The IB tradeoff is well-suited for bandwidth-limited edge inference and is a key design principle in our study for efficient communication. Wang et al. introduce the Informative Multi-Agent Communication (IMAC) method, which uses the information bottleneck principle to develop efficient communication protocols and scheduling for multi-agent reinforcement learning under limited bandwidth[45]. Shao et al. propose a task-oriented communication scheme for multi-device cooperative edge inference, optimizing local feature extraction and distributed feature encoding to minimize data redundancy and focus on task-relevant information, leveraging the information bottleneck principle and extending it to a distributed deterministic information bottleneck framework. However, these existing studies often neglect the need to prioritize different data from multiple cameras for various downstream tasks, such as considering ROIs. Moreover, most existing studies overlook the correlation between multiple cameras in multi-view scenarios. Shao et al. extract compact task-oriented representations based on the IB principle, but they overlook the fact that different tasks require varying levels of priority[16]. By leveraging these correlations and levels of priority, it is possible to further reduce data rates by minimizing duplicate information across different camera feeds and enhance inference performance at the same time.
@@ I-B2 Learning-Based Transmission Scheduling
Online learning in multi-agent deep reinforcement learning (MADRL) enhances multi-camera sensing under dynamic channels and overlapping ROIs. Effective transmission scheduling determines when and which agents communicate through binary vectors indicating allowed communications at specific time steps, forming a communication graph. Central transmission scheduling schemes use a globally shared policy to control communication. Kim et al. propose SchedNet, which uses a global scheduler to limit broadcasting agents and reduce communication overhead[46]. Du et al. introduce FlowComm, forming a directed graph for communication[47]. Liu et al. develop GA-Comm, using a two-stage attention network (G2ANet) to manage agent interactions[48]. Niu et al. present MAGIC, a framework using a directed communication graph for enhanced coordination[49]. In distributed transmission scheduling schemes, each agent individually determines whether to communicate, forming a graph structure. Liu et al. propose a framework for multi-agent collaborative perception, addressing communication group construction and decision-making for efficient bandwidth use, significantly reducing communication while maintaining performance[50]. Deep learning optimizes these systems by refining communication actions and schedules, transmitting only relevant information, and minimizing redundancy. However, existing multi-camera cooperative sensing algorithms do not effectively address the transmission scheduling problem, particularly under dynamic wireless channels and overlapping ROIs.
@@ I-C Our Contributions
  • We propose the PIB framework that prioritizes the shared data based on the signal-to-noise ratio (SNR) and camera coverage of the region of interest (RoI), reducing redundancy both spatially and temporally. This approach avoids the need for video reconstruction at edge servers and maintains low latency.
@@
  • To reduce communication costs in fluctuating links, we introduce a gate mechanism based on distributed online learning (DOL) to filter out unprofitable messages and efficiently select edge servers. We establish the asymptotic optimality of DOL by showing the sublinearity of its regrets.
@@ II System Model
As illustrated in Fig. 2, our system comprises a set of edge cameras, denoted as $\mathcal{K}=\{1,2,\ldots,K\}$, and a set of edge servers, denoted as $\mathcal{S}=\{1,2,\ldots,S\}$. Each camera has a specific Field of view (FoV), $\text{FoV}_{k}$, covering a subset of the monitored area. Our goal is to enable collaborative perception for pedestrian occupancy prediction under constrained channel capacity. The backbone of our system employs intermediate collaboration, where only encoded feature representations $Z$ are transmitted to a central aggregate server for decoding and inference. Other edge servers serve solely as relay nodes, as they do not perform decoding or final inference due to the network structure which does not support task offloading for these operations. The aggregate server is chosen to minimize overall system delay, while the relay-only role of other edge servers ensures streamlined data flow. The key notations are given in Table I for the ease of reading.
@@ III Problem Formulation
In this section, we establish the theoretical foundation for our PIB framework. We begin by detailing the IB analysis to determine the optimal balance between data compression and relevant information retention. Following this, we formulate the combinatorial multi-armed bandit (CMAB) problem to model the decision-making process of cameras in a distributed environment.
@@ III-A Prioritized Information Bottleneck Analysis
III-B Combinatorial Multi-armed Bandit (CMAB) Problem
@@
where $\Delta M_{k}=M_{\mathcal{K}\cup\{k\}}-M_{\mathcal{K}}$, and $\mathcal{K}$ is the set of cameras selected at time $t$. The computational cost of multi-camera fusion and inference at the edge server $s$ is denoted as $\Psi^{F}_{s}$. The computational cost of simply forwarding features from one edge server to another is denoted as $\Psi^{T}_{s}$. The remaining computing capacity of the $k$-th edge server is $\Psi^{R}_{k}$. $\mathcal{T}_{k,s}^{c\rightarrow e}$ denotes the latency of the transmission between the $k$-th camera and the $s$-th edge server, and $\mathcal{T}_{{s},{s}_{0}}^{e\rightarrow e}$ denotes the latency of the transmission between the aggregate node and the $s$-th edge server. Therefore, the CMAB problem can be formulated as:
@@
where (11a) ensures the number of selected cameras falls within the specified range and $|\mathcal{K}|=\sum_{k\in\mathcal{K}}{\mathcal{E}_{k,s}^{c\rightarrow e}}$, (11b) ensures that the remaining computing capacity of the aggregate server ($s_{0}$) chosen for inference is no less than the required capacity for fusion, (11c) ensures that each camera uses a unique transmission connection, (11d) ensures that the number of connections established by a single edge server does not exceed the maximum allowable connections, and (11e) ensures that the total latency for any transmission path is within the allowable time limit $\mathcal{T}^{U}$ for all edge servers ${s}\in\mathcal{S}$.